# Continual Distillation Learning: Knowledge Distillation in Prompt-based Continual Learning

### v3

@@ Abstract
We introduce the problem of continual distillation learning (CDL) in order to use knowledge distillation (KD) to improve prompt-based continual learning (CL) models. The CDL problem is valuable to study since the use of a larger vision transformer (ViT) leads to better performance in prompt-based continual learning. The distillation of knowledge from a large ViT to a small ViT can improve the inference efficiency for prompt-based CL models. We empirically found that existing KD methods such as logit distillation and feature distillation cannot effectively improve the student model in the CDL setup. To this end, we introduce a novel method named Knowledge Distillation based on Prompts (KDP), in which globally accessible prompts specifically designed for knowledge distillation are inserted into the frozen ViT backbone of the student model. We demonstrate that our KDP method effectively enhances the distillation performance in comparison to existing KD methods in the CDL setup. [^1] Project website https://irvlutd.github.io/CDL/
@@ 1 Introduction
Continual Learning (CL) [32] designs models that can continuously learn new tasks without forgetting previously learned tasks. For example, in class-incremental continual learning [26], the data of new classes arrive sequentially, and the model needs to learn to recognize these classes sequentially during training. In testing, the model will be tested on all the seen classes. Therefore, a good CL model should learn new classes without forgetting.
With recent advances in vision models, vision transformers (ViTs) [6] have demonstrated their advantages over convolutional neural networks (CNNs). Similarly, continual learning research has evolved from traditional CNN-based methods (e.g., using ResNet [11]) to the latest prompt-based CL methods that leverage ViTs as the backbone, which now represent the state-of-the-art CL models. Unlike traditional CL models where the backbone is trained, prompt-based CL methods such as L2P [35], DualPrompt [34] and CODA-Prompt [30], utilize a pre-trained ViT as the backbone while keeping it frozen during training. Instead, they optimize a prompt pool designed for continual learning. Using pre-trained ViT models and shifting learning to prompts, these methods achieve better performance compared to CNN-based continual learning approaches such as iCaRL [26] and LWF [19].
@@
We noticed that for prompt-based CL methods using ViTs, larger ViTs [6] achieve better performance. For example, ViT large is better than ViT base, and ViT base is better than ViT small, etc. Therefore, it is appealing to use large models in prompt-based CL. However, larger models introduce more computation during inference. To address this limitation, we propose a new research problem: Knowledge Distillation (KD) from large to small models in prompt-based continual learning. We name this problem Continual Distillation Learning (CDL), which aims to improve the performance of small models by using large models as teachers in continual learning.
@@
We particularly limit the scope of this work to prompt-based CL models, since we have experimentally verified that large models may not result in better performance for the traditional CNN-based CL models such as the iCaRL [26] model or the LWF [19] model (see Sec. 7 in supplementary material). These CNN-based models update the backbones during training. Deeper and larger CNNs tend to overfit new tasks, failing to fundamentally solve the catastrophic forgetting problem in continual learning.
@@
The CDL problem is different from the previous knowledge distillation setup. We illustrate the differences in Fig. 1. In the traditional KD setup [13], an offline dataset is used for knowledge distillation (Fig. 1(a)). A teacher model is first trained using the dataset, and then its knowledge is distilled to a student model using the same dataset. The setup is not the continual learning setup. The CDL problem that we study in this work is illustrated in Fig. 1(b). Given a new task in continual learning, a teacher model is first updated. Then a student model is updated based on the data of the new task and the teacher model. Consequently, a better student model can be trained by distilling knowledge from the teacher model. In general, this is a task-incremental learning process, where both the teacher and student models can only access the data for the current task at each step.
@@
To explore the new CDL problem, we first conducted an empirical study by applying existing knowledge distillation methods, i.e., logit distillation [13, 40], feature distillation [28, 4] and distillation token [31], to three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34], and CODA-Prompt [30]. For each model, we consider different teacher-student configurations. For example, we can use ViT-Large as the teacher model and ViT-Base as the student model. Since the ViT backbone is frozen in prompt-based CL, knowledge transfer from the teacher to the student can only be based on the prompt pool, where the selection of prompts relies on a key-query mechanism. However, for knowledge distillation, this mechanism is not effective. Since each task independently reselects prompt components from the prompt pool, the knowledge embedded within them from the teacher model cannot be transferred to the next task. Our experimental results also show that the performance improvements achieved by these knowledge distillation methods are not significant. Previous KD methods are not effective in the CDL setting.
@@
To address this limitation, we propose a novel method named Knowledge Distillation based on Prompts (KDP). KDP inserts globally accessible prompts, specifically designed for knowledge distillation, into the frozen ViT backbone of the student model. We name them KD prompts. These prompts are independent of the key-query mechanism of the prompt pool. They are not limited to specific tasks and can facilitate cross-task distillation. They serve as auxiliary information to guide the learning process of the student model. Experimental results demonstrate that this novel approach effectively addresses the CDL problem and enhances distillation performance compared to previous KD methods. Furthermore, we conducted an ablation study on the total number of KD prompts inserted into the backbone. The results show that the introduction of KD prompts into each ViT block significantly improves the alignment of the feature structures between layers of the teacher model and the student model, leading to better performance of the student model.
In summary, our contributions are as follows.
@@
  • We introduce the new problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning.
  • We conducted an empirical study by applying different knowledge distillation methods to three prompt-based CL approaches and identify the limitations of these distillation methods in the CDL setting.
@@
  • To address the unique challenges of the CDL problem, we propose a novel distillation method, KDP, which outperforms previous KD approaches.
@@ 2 Related Work
2.1 Continual Learning
The main purpose of continual learning is to build an intelligent system to solve the problem of catastrophic forgetting [24] in the case of incremental tasks. Different types of continual learning methods have been proposed in the literature [33]. For example, [26, 19] utilize function regularization to help the loss function. There are also weight regularization methods [17, 39, 22, 27, 29] that selectively constrain changes in network parameters and impose penalties on changes in each parameter based on its contribution. Architecture-based approaches [38, 15, 36, 7, 25] mainly focus on constructing a special model for continual learning. Replay-based approaches [26, 19, 14, 23, 1, 3, 2] use a memory buffer to store and replay old task data in learning the current task. Some recent works aim to address the issue of catastrophic forgetting without relying on rehearsal memory, which are referred as rehearsal-free methods [10, 8, 5]. The above methods are primarily based on CNN backbones, and a larger backbone model may not achieve better performance.
2.2 Prompt-based Continual Learning
Inspired by the use of prompts in natural language processing [21], prompt-based continual learning methods employ large vision transformer (ViT) models [6]. These methods freeze the pre-trained backbone and shift learning to prompts of the ViTs. For example, L2P [35] learns a prompt pool of key-prompt pairs, then selects the optimal prompt in the pool for a given input by matching the input with the keys in the pool. Based on this idea, DualPrompt [34] introduces the concept of using a general prompt and an expert prompt. CODA-Prompt [30] abandons the key-prompt pair selection idea. It uses a weighted sum of the prompt components to obtain the final prompt. CPrompt [9] introduces two main components during training: Classifier Consistency Learning (CCL) and Prompt Consistency Learning (PCL). In prompt-based methods, the size of the backbone matters. Larger backbone models achieve better performance. This motivates us to study the knowledge distillation problem for prompt-based continual learning methods, where we can distill knowledge from large models to small models in order to improve small models with fast inference time.
2.3 Knowledge Distillation
The purpose of knowledge distillation is to leverage larger models to assist smaller models to fully exploit their potential. The larger model, termed the teacher, aids in training the smaller model, termed the student. Traditional KD methods can be classified into logit distillation [13, 40] and feature distillation [20, 28, 4]. With recent rapid advancements in large models, new research has directed the focus of knowledge distillation towards vision transformers [31, 37]. However, KD for prompt-based continual learning models has not yet been explored, which is the focus of this work.
@@ 3.2 Prompt-based Methods
L2P: L2P [35] uses a prompt pool to encode information about tasks: $\mathcal{P}=\{\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M}\}$, where $\mathbf{P}_{i}\in\mathbb{R}^{L_{p}\times D}$ with token length $L_{p}$ and embedding size $D$, and $M$ is the total number of prompt components. To select prompts for different tasks, each prompt component is associated with a learnable key $\{(\mathbf{k}_{1},\mathbf{P}_{1}),(\mathbf{k}_{2},\mathbf{P}_{2}),\ldots,(\mathbf{k}_{M},\mathbf{P}_{M})\}$, where $\mathbf{k}_{i}\in\mathbb{R}^{D_{k}}$ with embedding size $D_{k}$ and $\mathcal{K}=\{\mathbf{k}_{i}\}_{i=1}^{M}$ denotes all the keys. Given an input image $\mathbf{x}\in\mathcal{X}$, a query function $q(\cdot)$ is used to encode the input image. The query $\mathbf{x}_{e}=q(\mathbf{x})\in\mathbb{R}^{D_{k}}$ then matches the key $\mathbf{k}_{i}$ with cosine similarity: $\gamma\left(q(\mathbf{x}),\mathbf{k}_{i}\right)$, where $\gamma(\cdot,\cdot)$ denotes the cosine similarity function. Then the top $K$ prompt components from the prompt pool $\mathcal{P}$ based on the establishment of key-value pairs are selected:
$\mathbf{P}_{\mathbf{x}}=\text{TopK}(\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M}),$ (1)
where TopK indicates the selection function based on cosine similarity. $\mathbf{P}_{\mathbf{x}}\in\mathbb{R}^{K\times L_{p}\times D}$ is the final set of prompt components extracted from the prompt pool to assist continual learning. The training loss function for a pair $(\mathbf{x},y)$ is defined as:
@@
CODA-Prompt: Instead of selecting the top $K$ key-value pairs in the prompt pool, CODA-Prompt [30] uses a weighted summation over the prompt components to compute the learnable prompt parameter for an input $\mathbf{x}$:
$\mathbf{P}_{\mathbf{x}}=\sum_{i=1}^{M}\alpha_{i}\mathbf{P}_{i},$ (3)
where $\alpha_{i}$ is the weight for the prompt component $\mathbf{P}_{i}$, which is computed using the cosine similarity of the query and key. Unlike L2P and DualPrompt, when calculating the similarity, CODA-Prompt creates a feature-selection attention scheme $\mathcal{A}=\{\mathbf{A}_{1},\ldots,\mathbf{A}_{M}\}$ for $M$ prompt components to process the query features, where $\mathbf{A}_{i}\in\mathbb{R}^{D_{k}},i=1,\ldots,M$. The weight is computed as $\alpha_{i}=\gamma\left(q(\mathbf{x})\odot\mathbf{A}_{i},\mathbf{k}_{i}\right)$, for $i=1,\ldots,M$. Finally, CODA-Prompt optimizes the prompt components $\mathcal{P}$, keys $\mathcal{K}$, attention $\mathcal{A}$ and the classifier. Additionally, orthogonality constraints are added to reduce interference between existing and new knowledge. The final loss function is:
@@
3.3 Prompting with Prefix-Tuning
After computing the prompt $\mathbf{P}\in\mathbb{R}^{L_{p}\times D}$ for an input sample, we need to insert this prompt into the ViT backbone for prompting. We use the prefix-tuning method. It divides the prompt $\mathbf{P}$ into two parts, $\mathbf{P}_{K}$, $\mathbf{P}_{V}\in\mathbb{R}^{L_{p}/2\times D}$, which are inserted into the key $\mathbf{h}_{K}$ and value $\mathbf{h}_{V}$ of the multi-head self-attention (MSA) layers, respectively:
@@ 3.3 Prompting with Prefix-Tuning
This method keeps the output sequence length the same as the input. In the CDL setup, since we need to insert prompts across multiple blocks, we employ the prefix-tuning method to avoid adding too many additional tokens.
@@ 4.1 Teacher-Student Model
As illustrated in Fig. 2, an input image $\mathbf{x}$ is first processed by the query function to generate a query, which is compared with keys $\mathcal{K}$ in the prompt pool. Different prompt-based continual learners select the prompts $\mathbf{P}^{\text{cl}}$ for continual learning through the Prompt Extractor, following Eq. (1) or Eq. (3). The prompts are divided into $\mathbf{P}^{\text{cl}}_{K}$ and $\mathbf{P}^{\text{cl}}_{V}$, which are inserted into the MSA layers across multiple blocks with prefix-tuning as in Eq. (5). This process is applied both the teacher and the student. Within this framework, we study and build different types of knowledge distillation in continual learning, including logit distillation and feature distillation.
@@
Feature distillation focuses on transferring the intermediate representations from a teacher model to a student model. In Fig. 2, the backbone ViT consists of multiple blocks, and we use the internal tokens outputted by each block as features. This allows the internal tokens of the student model to learn the information in the internal tokens of the teacher model. We built two CDL methods based on the handling of internal tokens: FitNets [28] and Review Knowledge Distillation (ReviewKD) [4].
Fig. 2 illustrates the logit distillation and the feature distillation in our continual distillation learning scenario.
Training of the teacher-student model. First, the images of the current task are trained on the teacher model. The teacher model is then used to help train the student model. During the student model training process, the teacher model remains frozen and is only used in the inference phase to provide soft labels (logit distillation) or internal tokens (feature distillation) to the student.
4.2 Limitations of Existing KD Methods in CDL

### v4

@@ Abstract
We introduce the problem of continual distillation learning (CDL) in order to use knowledge distillation (KD) to improve prompt-based continual learning (CL) models. The CDL problem is valuable to study since the use of a larger vision transformer (ViT) leads to better performance in prompt-based continual learning. The distillation of knowledge from a large ViT to a small ViT improves the inference efficiency for prompt-based CL models. We empirically found that existing KD methods such as logit distillation and feature distillation cannot effectively improve the student model in the CDL setup. To address this issue, we introduce a novel method named Knowledge Distillation based on Prompts (KDP), in which globally accessible prompts specifically designed for knowledge distillation are inserted into the frozen ViT backbone of the student model. We demonstrate that our KDP method effectively enhances the distillation performance in comparison to existing KD methods in the CDL setup.  [^1] Project website https://irvlutd.github.io/CDL/
@@ 1 Introduction
Continual Learning (CL) [32] designs models that can continuously learn new tasks without forgetting previously learned tasks. For example, in class-incremental continual learning [26], the data of new classes arrive sequentially, and the model needs to learn to recognize these classes sequentially during training. In testing, the model will be tested on all the seen classes. Therefore, a good CL model should learn new classes without forgetting. With recent advances in vision models, vision transformers (ViTs) [6] have demonstrated their advantages over convolutional neural networks (CNNs). Similarly, continual learning research has evolved from traditional CNN-based methods (e.g., using ResNet [11]) to the latest prompt-based CL methods that leverage ViTs as the backbone, which now represent the state-of-the-art CL models. Unlike traditional CL models where the backbone is trained, prompt-based CL methods such as L2P [35], DualPrompt [34] and CODA-Prompt [30], utilize a pre-trained ViT as the backbone while keeping it frozen during training. Instead, they optimize a prompt pool designed for continual learning. Using pre-trained ViT models and shifting learning to prompts, these methods achieve better performance compared to CNN-based continual learning approaches such as iCaRL [26] and LWF [19].
@@
We noticed that for prompt-based CL methods using ViTs, larger ViTs [6] achieve better performance. For example, ViT large is better than ViT base, and ViT base is better than ViT small, etc. Therefore, it is appealing to use large models in prompt-based CL. However, larger models introduce more computation during inference. To address this limitation, we propose a new research problem: Knowledge Distillation (KD) from large to small models in rehearsal-free prompt-based continual learning. We name this problem Continual Distillation Learning (CDL), which aims to improve the performance of small models by using large models as teachers in continual learning.
@@
We particularly limit the scope of this work to prompt-based CL models, since we have experimentally verified that large models may not result in better performance for the traditional CNN-based CL models such as iCaRL [26] or LWF [19] (see Sec. A in Appendix). These CNN-based models update the backbones during training. Deeper and larger CNNs tend to overfit new tasks, failing to fundamentally solve the catastrophic forgetting problem.
@@
The CDL problem is different from the previous knowledge distillation setup as illustrated in Fig. 1. In the traditional KD setup [13], an offline dataset is used for knowledge distillation (Fig. 1(a)). A teacher model is first trained using the dataset, and then its knowledge is distilled to a student model using the same dataset. The setup is not the continual learning setup. The CDL problem that we study in this work is illustrated in Fig. 1(b). Given a new task in continual learning, a teacher model is first updated. Then a student model is updated based on the data of the new task and the teacher model. Consequently, a better student model can be trained by distilling knowledge from the teacher model. In general, this is a task-incremental learning process, where both the teacher and student models can only access the data for the current task at each step. It is worth noting that some traditional continual learning methods also involve knowledge distillation, but in those cases, the teacher is typically the model trained on previous tasks under the same architecture. In contrast, our CDL framework is based on a fundamentally different perspective: it investigates the process where, in each task, a student model learns from a teacher model with a larger ViT backbone. This distillation process is expected to continue throughout the entire task-incremental learning procedure.
@@
When applying existing knowledge distillation methods, i.e., logit distillation [13, 40], feature distillation [28, 4], to the CDL problem, we found that the results are not satisfactory. Since the ViT backbone is frozen in prompt-based CL, knowledge transfer from the teacher to the student can only be based on the prompt pool. However, for knowledge distillation, this mechanism is not effective. Since each task independently reselects prompt components from the prompt pool, the knowledge embedded within them from the teacher model cannot be transferred to the next task. We refer to this phenomenon as distillation information forgetting.
@@
To address this problem, we propose a novel method named Knowledge Distillation based on Prompts (KDP). KDP inserts globally accessible prompts, specifically designed for knowledge distillation, into the frozen ViT backbone of the student model. We name them KD prompts. These prompts are independent of the key-query mechanism of the prompt pool. They are not limited to specific tasks and can facilitate cross-task distillation. They serve as auxiliary information to guide the learning process of the student model. To enhance generalization, we conducted experiments on three representative prompt-based CL methods: L2P, DualPrompt, and CODA-Prompt. The results demonstrate that our KDP method effectively addresses this issue and achieves state-of-the-art performance compared to other distillation approaches. In summary, our contributions are as follows:
@@
  • We introduce the new problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning and conduct an empirical study by applying different knowledge distillation methods to different prompt-based CL approaches.
  • We are the first to discover and propose the distillation information forgetting problem presented in existing distillation methods and provide an analysis of its underlying causes.
@@
  • To address this problem, we propose a novel distillation method, KDP, which outperforms previous KD approaches in continual distillation learning.
@@ 2 Related Work
Continual Learning. The main purpose of continual learning is to build an intelligent system to solve the problem of catastrophic forgetting [24] in the case of incremental tasks. Different types of continual learning methods have been proposed in the literature [33]. For example, [26, 19] utilize function regularization to help the loss function. There are also weight regularization methods [17, 39, 22, 27, 29] that selectively constrain changes in network parameters and impose penalties on changes in each parameter based on its contribution. Architecture-based approaches [38, 15, 36, 7, 25] mainly focus on constructing a special model for continual learning. Replay-based approaches [26, 19, 14, 23, 1, 3, 2] use a memory buffer to store and replay old task data in learning the current task. Some recent works aim to address the issue of catastrophic forgetting without relying on rehearsal memory, which are referred as rehearsal-free methods [10, 8, 5]. The above methods are primarily based on CNN backbones, and a larger backbone model may not achieve better performance.
Prompt-based Continual Learning. Inspired by the use of prompts in natural language processing [21], prompt-based continual learning methods employ large vision transformer (ViT) models [6]. These methods freeze the pre-trained backbone and shift learning to prompts of the ViTs. For example, L2P [35] learns a prompt pool of key-prompt pairs, then selects the optimal prompt in the pool for a given input by matching the input with the keys in the pool. Based on this idea, DualPrompt [34] introduces the concept of using a general prompt and an expert prompt. CODA-Prompt [30] abandons the key-prompt pair selection idea. It uses a weighted sum of the prompt components to obtain the final prompt. CPrompt [9] introduces two main components during training: Classifier Consistency Learning (CCL) and Prompt Consistency Learning (PCL). In prompt-based methods, larger backbone models achieve better performance. This motivates us to study the knowledge distillation problem for prompt-based continual learning methods, where we can distill knowledge from large models to small models in order to improve small models with fast inference time.
Knowledge Distillation. The purpose of knowledge distillation is to leverage larger models to improve smaller models. The larger model, termed the teacher, aids in training the smaller model, termed the student. Traditional KD methods can be classified into logit distillation [13, 40] and feature distillation [20, 28, 4]. With recent rapid advancements in large models, new research has directed the focus of knowledge distillation towards vision transformers [31, 37]. However, KD for prompt-based continual learning models has not yet been explored, which is the focus of this work.
@@ 3 Prerequisites
L2P: L2P [35] uses a prompt pool to encode information about tasks: $\mathcal{P}=\{\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M}\}$, where $\mathbf{P}_{i}\in\mathbb{R}^{L_{p}\times D}$ with token length $L_{p}$ and embedding size $D$, and $M$ is the total number of prompt components. To select prompts for different tasks, each prompt component is associated with a learnable key $\{(\mathbf{k}_{1},\mathbf{P}_{1}),(\mathbf{k}_{2},\mathbf{P}_{2}),\ldots,(\mathbf{k}_{M},\mathbf{P}_{M})\}$, where $\mathbf{k}_{i}\in\mathbb{R}^{D_{k}}$ with embedding size $D_{k}$ and $\mathcal{K}=\{\mathbf{k}_{i}\}_{i=1}^{M}$ denotes all the keys. Given an input image $\mathbf{x}\in\mathcal{X}$, a query function $q(\cdot)$ is used to encode the input image. The query $\mathbf{x}_{e}=q(\mathbf{x})\in\mathbb{R}^{D_{k}}$ then matches the key $\mathbf{k}_{i}$ with cosine similarity: $\gamma\left(q(\mathbf{x}),\mathbf{k}_{i}\right)$, where $\gamma(\cdot,\cdot)$ denotes the cosine similarity function. Then the top $K$ prompt components from the prompt pool $\mathcal{P}$ based on the establishment of key-value pairs are selected: $\mathbf{P}_{\mathbf{x}}=\text{TopK}(\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M})$, where TopK indicates the selection function based on cosine similarity. $\mathbf{P}_{\mathbf{x}}\in\mathbb{R}^{K\times L_{p}\times D}$ is the final set of prompt components extracted from the prompt pool to assist continual learning. The training loss function for a pair $(\mathbf{x},y)$ is defined as:
@@
CODA-Prompt: Instead of selecting the top $K$ key-value pairs in the prompt pool, CODA-Prompt [30] uses a weighted summation over the prompt components to compute the learnable prompt parameter for an input $\mathbf{x}$: $\mathbf{P}_{\mathbf{x}}=\sum_{i=1}^{M}\alpha_{i}\mathbf{P}_{i}$, where $\alpha_{i}$ is the weight for the prompt component $\mathbf{P}_{i}$, which is computed using the cosine similarity of the query and key. Unlike L2P and DualPrompt, when calculating the similarity, CODA-Prompt creates a feature-selection attention scheme $\mathcal{A}=\{\mathbf{A}_{1},\ldots,\mathbf{A}_{M}\}$ for $M$ prompt components to process the query features, where $\mathbf{A}_{i}\in\mathbb{R}^{D_{k}},i=1,\ldots,M$. The weight is computed as $\alpha_{i}=\gamma\left(q(\mathbf{x})\odot\mathbf{A}_{i},\mathbf{k}_{i}\right)$, for $i=1,\ldots,M$. Finally, CODA-Prompt optimizes the prompt components $\mathcal{P}$, keys $\mathcal{K}$, attention $\mathcal{A}$ and the classifier. Additionally, orthogonality constraints are added to reduce interference between existing and new knowledge. The final loss function is:
@@
Prompting with Prefix-Tuning: After computing the prompt $\mathbf{P}\in\mathbb{R}^{L_{p}\times D}$ for an input sample, we need to insert this prompt into the ViT backbone for prompting. We use the prefix-tuning method. It divides the prompt $\mathbf{P}$ into two parts, $\mathbf{P}_{K}$, $\mathbf{P}_{V}\in\mathbb{R}^{L_{p}/2\times D}$, which are inserted into the key $\mathbf{h}_{K}$ and value $\mathbf{h}_{V}$ of the multi-head self-attention (MSA) layers, respectively:
@@
This method keeps the output sequence length the same as the input. In CDL, we employ the prefix-tuning method to avoid adding too many additional tokens.
@@ 4.1 Teacher-Student Model
As illustrated in Fig. 2, an input image $\mathbf{x}$ is first processed by the query function to generate a query, which is compared with keys $\mathcal{K}$ in the prompt pool. Different prompt-based continual learners select the prompts $\mathbf{P}^{\text{cl}}$ for continual learning through the Prompt Extractor, such as using TopK or weighted combination. The prompts are divided into $\mathbf{P}^{\text{cl}}_{K}$ and $\mathbf{P}^{\text{cl}}_{V}$, which are inserted into the MSA layers across multiple blocks with prefix-tuning as in Eq. (3). This process is applied both the teacher and the student. Within this framework, we study and build different types of knowledge distillation in continual learning, including logit distillation and feature distillation.
@@
Feature distillation focuses on transferring the intermediate representations from a teacher model to a student model. In Fig. 2, the backbone ViT consists of multiple blocks, and we use the internal tokens outputted by each block as features. This allows the internal tokens of the student model to learn the information in the internal tokens of the teacher model. We built two CDL methods based on the handling of internal tokens: FitNets [28] and Review Knowledge Distillation (ReviewKD) [4]. Fig. 2 illustrates the logit distillation and the feature distillation in our CDL scenario.
Training of the teacher-student model. First, the images of the current task are trained on the teacher model. The teacher model is then used to help train the student model. During the student model training process, the teacher model remains frozen and is only used in the training phase to provide soft labels (logit distillation) or internal tokens (feature distillation) to the student.
4.2 Distillation Information Forgetting
Figure 3: (a) A student model may select different prompts for each new task, where teacher information is lost during distillation. (b) To address this limitation, one solution is to unfreeze part of the ViT, which results in a severe forgetting problem. (c) Our proposed solution is to introduce KD prompts, treating the ViT backbone together with the KD prompts as an overall trainable structure.